# PDDFormer: Pairwise Distance Distribution Graph Transformer for Crystal Material Property Prediction

### v4

@@ 3 Related work
Finite crystal graph representation. CGCNN Xie and Grossman (2018) predicts material properties by learning the connections between atoms in crystals by representing crystal structures as finite multi-edge crystal graphs. Building on the construction of multi-edge crystal graphs, MegNet Chen et al. (2019) introduced global state attributes into graph networks, while GATGNN Louis et al. (2020) utilized multiple graph attention layers (GAT) to learn the properties of local neighborhoods and employed global attention layers to weight global atomic features. ALIGNN Choudhary and DeCost (2021) and M3GNet Chen and Ong (2022) incorporated angular information into the message-passing process to generate richer and more discriminative representations. CrysMMNet Das et al. (2023) adopted a multimodal framework, integrating graph and text representations to produce joint multimodal representations of crystalline materials. CrysDiff Song et al. (2024) is a pretraining-finetuning framework based on diffusion models. However, the aforementioned methods represent crystals as finite graph structures, failing to capture the periodicity of infinite crystals effectively.
@@
Periodic representation of crystals. Recently, Matformer Yan et al. (2022) encoded periodic patterns by adding self-connecting edges to atoms based on lattice parameters, directly using lattice parameters to encode periodic structures under ideal conditions. PotNet Lin et al. (2023) considered the infinite summation of interatomic interactions. Crystalformer Taniai et al. (2024) performed infinite summations of interatomic potentials through infinitely connected attention while also utilizing lattice parameters. ComFormer Yan et al. (2024a) constructed cell parameters by adding self-connecting edges to atoms and their copies in three different directions to encode periodic patterns, employing equivariant vector representations and invariant geometric descriptors of Euclidean distances and angles to represent the geometric information of crystals. GMTNet Yan et al. (2024b) aims to predict the tensor properties of crystalline materials while satisfying O(3) group equivariance and the symmetry of crystal space groups to ensure the accuracy and consistency of tensor predictions. However, while these methods achieve complete crystal graph representations, the crystal structures they represent rely on non-invariants or discontinuous invariants, such as lattice parameters, symmetry, and space groups, failing to address the issue of crystal data fuzziness.
@@
Continuity and complete representations for crystals. Addressing the continuity and completeness of crystal representations is a critical issue. Recent advancements in AMD Wang et al. (2022) and PDD Widdowson and Kurlin (2022) have developed matrix forms that are both complete and continuous. However, in practical applications, using these matrix representations as inputs for predicting crystal properties without compromising continuity and completeness is challenging. The AMD and PDD representations are designed for stable crystal structures and do not account for atomic types. Their completeness relies on the assumption that no two crystals with identical structures differ solely by atomic type, which is only feasible for stable structures. Additionally, to achieve completeness, a sufficiently large number of neighbors $k$ must be predetermined for any test crystal. Typically, hundreds of neighbors are required to distinguish all periodic structures in the Cambridge Structural Database. Directly modeling PDD as edge information is impractical and costly in real-world applications Balasingham et al. (2022).
@@
Figure 2: Schematic diagram of the selected neighbors in PDD in 2D. The edges in Figure a show the neighbor selection for atom $i$ in WPDD, represented by the red lines $d_{1},d_{2},d_{3}$. They are ordered by Euclidean distance as $d_{1}<d_{2}<d_{3}$. By comparing Figures (b) and (c), we can see that we construct the unit cell centered around each atom and select neighbors, rather than being limited to the unit cell where the atoms are located.
@@ 4 PDDFormer
In this section, we first propose two variants of PDD, namely WPDD and UPDD, and then incorporate them into crystal graph construction. We finally present the PDDformer framework.
4.1 WPDD
@@
$\begin{array}[]{@{}c@{}c@{}}WPDD=\left(W,PDD\right)=\\ {\textstyle\bigcup_{i=1}^{n}}\left(\frac{t(x_{i})}{\sum_{j=1}^{n}t(x_{j})},{\textstyle\bigcup_{j=1}^{k}}\sqrt{\left(p_{i}-p_{j}\right)^{2}}\right)\end{array}$ (2)
[…]
4.2 UPDD
@@
$\begin{array}[]{@{}c@{}c@{}}\mathcal{D}_{i}=\left\{{\textstyle\bigcup_{j=1}^{n}}\sqrt{\left(\overline{p_{i}}-\overline{p_{j}}\right)^{2}}\mid i,j\in\mathcal{Z}\right\},\\ \left(\overline{p_{i}}-\overline{p_{j}}\right)^{2}=\left(\overline{x_{i}}-\overline{x_{j}}\right)^{2}+\left(\overline{y_{i}}-\overline{y_{j}}\right)^{2}+\left(\overline{z_{i}}-\overline{z_{j}}\right)^{2},\\ UPDD=\left\{\bigcup_{i=1}^{n}\frac{1}{\mathcal{D}_{i}}\mid i\in\mathcal{Z},\mathcal{D}_{i}\neq 0\right\}\end{array}$ (3)
@@
As shown in Figure 2, the selection is not based on Euclidean distances, but rather on choosing atoms within the reconstructed unit cell for construction. The dimension of our UPDD is determined by the atoms in the unit cell and does not require consideration of the neighbor count, k, across different datasets, making it more generalizable. This UPDD covers unit cell structures with a larger number of atoms while also ensuring that unit cell structures with fewer atoms are not disturbed by excessive neighbor information. It also reduces resource consumption. Due to this crystal-specific treatment, the UPDD dimensions of different crystal structures may not match, so dimension alignment is required before feeding into the neural network.
@@ 4.3 Crystal graph construction
By introducing PDD, we constructed a complete and continuous multi-edge crystal graph. In the graph, each node represents an atom $i$ and all its infinite duplicates in 3D space, with positions $\{\hat{p}_{i}|\hat{p}_{i}=p_{i}+k_{1}l_{1}+k_{2}l_{2}+k_{3}l_{3},k_{1},k_{2},k_{3}\in\mathcal{Z}\}$, and node features $x_{i}$. An edge is established from node $j$ to node $i$ when the Euclidean distance $\left|e_{j^{\prime}i}\right|^{2}$ between a duplicate of $j$ and $i$ satisfies $\left|e_{j^{\prime}i}\right|^{2}=\left|p_{j}+k^{\prime}_{1}l_{1}+k^{\prime}_{2}l_{2}+k^{\prime}_{3}l_{3}-p_{i}\right|^{2}\leq r$, where $r\in\mathbb{R}$ is the cutoff radius. We select the nearest $t$ edges within the cutoff radius, each with a corresponding edge feature $\left|e_{j^{\prime}i}\right|^{2}$. Since WPDD requires a large number of neighbors to be predefined, representing this neighbor information as edge features is neither practical nor cost-effective in real-world applications. Therefore, we retain its matrix form and incorporate it into the construction of the multi-edge crystal graph as a way to reflect the global information of the crystal structure. After passing through the Embedding Block in Section 4.6, UPDD is aligned in dimensions and transformed into matrix form data. Formally, we represent the constructed crystal graph as $\mathcal{G}=(\mathcal{X},\mathcal{XI},\mathcal{E},PDD)$. Therein, $x_{i}\in\mathcal{X}$ is the feature vector of the atom $i$, $e_{ij}^{h}\in\mathcal{E}$ is the feature vector of the $h$-th edge between nodes $i$ and $j$, and we denote $\mathcal{XI}$ as the indices of the nodes $i$ and $j$ that form the edge. Sections 4.4 and 4.5 are our proofs of the continuity and geometric completeness of PDD crystal graphs.
4.4 Continuity of Proposed Crystal Graphs
@@
The continuity of the constructed crystal graph $\mathcal{G}\left(\mathcal{S}\right)$ under perturbations of the crystal structure $\mathcal{S}$ will be measured using the EMD Rubner et al. (2000), which applies to crystal graphs of any size. Definition 6 applies to any crystal graph $\mathcal{G}\left(\mathcal{S}\right)=\left(\left[w_{1}\left(S\right),R_{1}\left(S\right)\right],\dots,\left[w_{1}\left(S\right),R_{1}\left(S\right)\right]\right)$, where $\left[w_{i}\left(S\right),R_{i}\left(S\right)\right]$ represent the information extracted based on atom $i$ in the unit cell. $R_{i}\left(S\right)=R_{i}\left(\mathcal{S}_{\mathcal{X}},\mathcal{S}_{\mathcal{XI}},\mathcal{S}_{\mathcal{E}},\mathcal{S}_{PDD}\right)$ includes atomic information, neighbor information used in constructing the multi-edge crystal graph, and the PDD invariants of the crystal structure $\mathcal{S}$, with weights $w_{i}\in\left(0,1\right]$ satisfying the normalization condition ${\textstyle\sum_{i=1}^{n}}w_{i}\left(\mathcal{S}\right)=1$.
@@
Subsequently, we only consider the case where the weighted vector $\left[w_{i},R_{i}\right]$ corresponds to the $i$-th row of PDD$\left(S;k\right)$. Here, $n$ denotes the number of rows in PDD$\left(S;k\right)$. The size of each row $R_{i}\left(S\right)$ should be independent of $\mathcal{S}$ and depend solely on the number of neighbors $k$ in PDD$\left(S;k\right)$. For any vectors $R_{i}=\left(r_{i1},\dots,r_{ik}\right)$ and $R_{j}=\left(r_{j1},\dots,r_{jk}\right)$ of length $k$, we use the $L_{\infty}$- distance $\left|R_{i}-R_{j}\right|_{\infty}=\max_{l=0,\dots k}\left|r_{il}-r_{jl}\right|_{\infty}$.
Proposition 1. The WPDD and UPDD multi-edge crystal graph is continuous.
@@
Proof: For any $k\geq 1$, if the periodic crystal $\mathcal{S},\mathcal{Q}\in\mathbb{R}^{n}$ satisfy $d_{B}\left(\mathcal{S},\mathcal{Q}\right)<r\left(\mathcal{S}\right)$, then we have:$EMD\left(\mathcal{G}\left(\mathcal{S}\right),\mathcal{G}\left(\mathcal{Q}\right)\right)=EMD\left(\left(\mathcal{S}_{\mathcal{X}},\mathcal{S}_{\mathcal{XI}},\mathcal{S}_{\mathcal{E}},\mathcal{S}_{PDD}\right),\left(\mathcal{Q}_{\mathcal{X}},\mathcal{Q}_{\mathcal{XI}},\mathcal{Q}_{\mathcal{E}},\mathcal{Q}_{PDD}\right)\right)=EMD\left(\left(\mathcal{S}_{\mathcal{X}},\mathcal{Q}_{\mathcal{X}}\right)\right)+EMD\left(\left(\mathcal{S}_{\mathcal{XI}},\mathcal{Q}_{\mathcal{XI}}\right)\right)+EMD\left(\left(\mathcal{S}_{\mathcal{E}},\mathcal{Q}_{\mathcal{E}}\right)\right)+EMD\left(\left(\mathcal{S}_{PDD},\mathcal{Q}_{PDD}\right)\right)$. Since disturbances only change the positions of atoms and do not alter their types, therefore $EMD\left(\left(\mathcal{S}_{\mathcal{X}},\mathcal{Q}_{\mathcal{X}}\right)\right)=0$ and $EMD\left(\left(\mathcal{S}_{\mathcal{XI}},\mathcal{Q}_{\mathcal{XI}}\right)\right)=0$. So, we obtain $EMD\left(\mathcal{G}\left(\mathcal{S}\right),\mathcal{G}\left(\mathcal{Q}\right)\right)=EMD\left(\left(\mathcal{S}_{\mathcal{E}},\mathcal{Q}_{\mathcal{E}}\right)\right)+EMD\left(\left(\mathcal{S}_{PDD},\mathcal{Q}_{PDD}\right)\right)\leq 2d_{B}\left(\mathcal{S},\mathcal{Q}\right)$.
@@
The bottleneck distance $d_{B}\left(\mathcal{S},\mathcal{Q}\right)<r\left(\mathcal{S}\right)$ is defined as: $d_{B}\left(\mathcal{S},\mathcal{Q}\right)=\inf_{g:\mathcal{S}\to\mathcal{Q}}\sup_{p\in\mathcal{S}}\left|p-g\left(p\right)\right|$ and the envelope radius $r\left(\mathcal{S}\right)$ is the minimum half-distance between any two points in $r\left(\mathcal{S}\right)$. In other words, $r\left(\mathcal{S}\right)$ is the maximum radius of non-overlapping open balls centered at all points in $\mathcal{S}$. This implies that any small perturbation in atomic positions under the $d_{B}$ Carstens et al. (1999) will lead to minor changes in the distribution of distances between points in the EMD.
@@
Since the EMD between the constructed crystal graphs only relates to Euclidean distance. Euclidean distance itself is continuous, Theorem 1 extends the following fact: for a unit cell structure with two atoms, when the number of neighbors $k=1$, if we perturb at most two points by $\epsilon$, the change in distance between the two points will be at most $2\epsilon$. Extending this to $n$ atomic points with $k$ neighbors, if we perturb at most $n$ points by $\epsilon$, the change in distance between $n$ points will be at most $2nk\epsilon$. This aligns with Definition 5, hence the constructed WPDD and UPDD multi-edge crystal graph is continuous.
Figure 3: Illustrations of different unit cell and lattice representations of the same crystal structure. The blue area in the figure represents possible unit cell structures. Figure (a) shows several possible choices among the infinitely many unit cells for the same crystal structure in the undisturbed case. Figure (b) illustrates that for almost any perturbation, the symmetry group and any reduced unit cell (with minimal volume) will undergo discontinuous changes.
4.5 Geometric completeness of proposed crystal graphs

### v5

@@ 3 Related Work
CGCNN xie2018crystal represents crystal structures as finite multi-edge crystal graphs to model crystal structures and predict material properties. Building on the construction of multi-edge crystal graphs, MegNet chen2019graph introduced global state attributes into graph networks, while GATGNN louis2020graph utilized multiple graph attention layers (GAT) to learn the properties of local neighborhoods and employed global attention layers to weight global atomic features. ALIGNN choudhary2021atomistic and M3GNet chen2022universal incorporated angular information into the message-passing process to generate richer and more discriminative representations.  das2022crysxpp proposed an interpretable deep property predictor called CrysXPP. CrysMMNet das2023crysmmnet adopted a multimodal framework, integrating graph and text representations to produce joint multimodal representations of crystalline materials.  das2023crysgnn proposed CrysGNN, a pretraining framework leveraging unlabeled data, while CrysDiff song2024diffusion introduced a diffusion model-based pretraining–fine-tuning framework. However, the methods above represent crystals as finite graph structures, which fail to effectively capture the periodicity of infinite crystals.
Periodic Representation of Crystals
@@
Recently, Matformer yan2022periodic encoded periodic patterns by adding self-connecting edges to atoms based on lattice parameters. PotNet lin2023efficient considered the infinite summation of interatomic interactions. Crystalformer taniai2024crystalformer performed infinite summations of interatomic potentials through infinitely connected attention while also utilizing lattice parameters. ComFormer yan2024complete constructed cell parameters by adding self-connecting edges to atoms and their copies in three different directions to encode periodic patterns, employing equivariant vector representations and invariant geometric descriptors of Euclidean distances and angles to represent the geometric information of crystals. GMTNet yan2024space aims to predict the tensor properties of crystalline materials while satisfying O(3) group equivariance and the symmetry of crystal space groups. However, the crystal structures they represent rely on non-invariant or discontinuous invariants, failing to resolve the issue of crystal data fuzziness.
Continuity and Complete Representations for Crystals
@@
Addressing the continuity and completeness of crystal representations is a critical issue. Recent advancements in AMD wang2022approximately and PDD widdowson2022resolving have developed matrix forms that are both complete and continuous. However, in practical applications, using these matrix representations as inputs for predicting crystal properties without compromising continuity and completeness is challenging. The AMD and PDD representations are designed to distinguish stable crystal structures and do not consider atomic types; their completeness assumption only holds for stable structures. Additionally, to achieve completeness, a sufficiently large number of neighbors $k$ must be predetermined for any test crystal. Directly modeling PDD as edge information is impractical and costly in real-world applications balasingham2022compact.
@@
Figure 2: Schematic diagram of the selected neighbors in PDD. Figure (a) represents the 3D unit cell structure. The edges in Figure (b) show the neighbor selection for atom $i$ in WPDD. By comparing Figures (c) and (d), we can see that we construct the unit cell centered around each atom and select neighbors, rather than being limited to the unit cell where the atoms are located.
@@ 4 PDDFormer
In this section, we propose two PDD variants, namely WPDD and UPDD, and then incorporate them into crystal graph construction. We finally present the PDDformer framework.
4.1 Atom-Weighted PDD (WPDD)
Since the PDD representation is designed for stable crystal structures and does not consider atomic types, it is not suitable for predicting crystal material properties. To account for the influence of atomic types, for a given crystal structure $\mathcal{S}=\mathcal{U}+\mathcal{L}$, where each atom $x_{i}\in\mathcal{U}$ is labeled with the atomic mass $t(x_{i})$ corresponding to it, the final weight for each row is $\mathcal{W}=[w_{1},\dots,w_{n}]^{T}$, where $w_{i}=\frac{t(x_{i})}{\sum_{u=1}^{n}t(x_{u})}$. By concatenating $\mathcal{W}$ with $\mathcal{PDD}\in\mathbb{R}^{n\times k}$, an atomic-mass-weighted $\mathcal{WPDD}\left(S;k\right)\in\mathbb{R}^{n\times\left(k+1\right)}$ is formed, represented by the following equation:
@@ 4.1 Atom-Weighted PDD (WPDD)
Herein, $n$ represents the number of atoms in the unit cell, and $p_{i}$ and $p_{j}$ denote the spatial coordinates of an atom $i$ and its neighbor $j$, respectively, and $k$ is the number of nearest neighbors selected when constructing the PDD, sorted in ascending order of Euclidean distance as $d_{i1}\leq\dots\leq d_{ik}$, as shown in Figure 2(b). WPDD is equivalent to the PDD of the crystal structure $\mathcal{S}$, except that the rows are not grouped as in the original version. This prevents the loss of atomic information when two primitive points have the same $k$-nearest neighbor distances but correspond to different atomic types. Therefore, $\mathcal{WPDD}\in\mathbb{R}^{n\times\left(k+1\right)}$, where $n$ is the number of atoms in the constructed graph.
4.2 Unit-cell PDD (UPDD)
@@
Since the interaction energy between an atom and its neighboring atoms is usually inversely proportional to the distance, we take the reciprocal feature of the distance after removing zeros.
As shown in Figure 2(d), the selection is not based on Euclidean distances but on choosing atoms within the reconstructed unit cell for construction. The atoms in the unit cell determine the dimension of our UPDD and do not require consideration of the neighbor count, k, across different datasets, making it more generalizable. This UPDD covers unit cell structures with a larger number of atoms while ensuring that unit cell structures with fewer atoms are not disturbed by excessive neighbor information. It also reduces resource consumption. Due to this crystal-specific treatment, the UPDD dimensions of different crystal structures may not match, so dimension alignment is required before feeding them into the neural network.
@@ 4.3 Crystal Graph Construction
By introducing PDD, we constructed a general complete and continuous multi-edge crystal graph. In the graph, node features are $x_{i}$. An edge is established from node $j$ to node $i$ when the Euclidean distance $\left|e_{j^{\prime}i}\right|^{2}$ between a duplicate of $j$ and $i$ satisfies $\left|e_{j^{\prime}i}\right|^{2}=\left|p_{j}+k^{\prime}_{1}l_{1}+k^{\prime}_{2}l_{2}+k^{\prime}_{3}l_{3}-p_{i}\right|^{2}\leq r$, where $r\in\mathbb{R}$ is the cutoff radius. Next, we construct a PDD row for each atom. Since directly representing PDD as edge features is impractical, we retain its matrix form and incorporate it into the construction of the multi-edge crystal graph to reflect the global information of the crystal structure. Therefore, we represent the constructed crystal graph as $\mathcal{G}=(\mathcal{X},\mathcal{XI},\mathcal{E},\mathcal{PDD})$. Therein, $x_{i}\in\mathcal{X}$ is the feature vector of the atom $i$, $e_{ij}^{h}\in\mathcal{E}$ is the feature vector of the $h$-th edge between nodes $i$ and $j$, and we denote $\mathcal{XI}$ as the indices of the nodes $i$ and $j$ that form the edge. Sections 4.4 and 4.5 are our proofs of the continuity and general geometric completeness of PDD crystal graphs.
Figure 3: Architecture Overview. PDDFormer accepts an input crystal structure $S$. During the prediction process, it first undergoes a graph construction step to generate a continuous and general complete crystal graph structure, followed by an embedding block, then multiple blocks of node-wise Transformer and PDD Message Passing, and finally, an output block.
4.4 The Continuity of Proposed Crystal Graphs
@@
The continuity of the constructed crystal graph $\mathcal{G}\left(\mathcal{S}\right)$ under perturbations of the crystal structure $\mathcal{S}$ will be measured using the EMD rubner2000earth, which applies to crystal graphs of any size. Definition 6 applies to any crystal graph $\mathcal{G}\left(\mathcal{S}\right)=\left(\left[w_{1}\left(S\right),R_{1}\left(S\right)\right],\dots,\left[w_{n}\left(S\right),R_{n}\left(S\right)\right]\right)$, where $\left[w_{i}\left(S\right),R_{i}\left(S\right)\right]$ represent the information extracted based on atom $i$ in the unit cell. $R_{i}\left(S\right)=R_{i}\left(\mathcal{S}_{\mathcal{X}},\mathcal{S}_{\mathcal{XI}},\mathcal{S}_{\mathcal{E}},\mathcal{S}_{\mathcal{PDD}}\right)$ includes atomic information, neighbor information used in constructing the multi-edge crystal graph, and the PDD invariants of the crystal structure $\mathcal{S}$, with weights $w_{i}\in\left(0,1\right]$ satisfying the normalization condition ${\textstyle\sum_{i=1}^{n}}w_{i}\left(\mathcal{S}\right)=1$.
@@
Subsequently, we only consider the case where the weighted vector $\left[w_{i},R_{i}\right]$ corresponds to the $i$-th row of $\mathcal{PDD}\left(S;k\right)$. Here, $n$ denotes the number of rows in $\mathcal{PDD}\left(S;k\right)$. The size of each row $R_{i}\left(S\right)$ should be independent of $\mathcal{S}$ and depend solely on the number of neighbors $k$ in $\mathcal{PDD}\left(S;k\right)$. For any vectors $R_{i}=\left(r_{i1},\dots,r_{ik}\right)$ and $R_{j}=\left(r_{j1},\dots,r_{jk}\right)$ of length $k$, we use the $L_{\infty}$- distance $\left|R_{i}-R_{j}\right|_{\infty}=\max_{l=0,\dots k}\left|r_{il}-r_{jl}\right|_{\infty}$.
Proposition 1.
The WPDD and UPDD multi-edge crystal graph is continuous.
Proof.
For any neighbor count $k\geq 1$, if the periodic crystal $\mathcal{S},\mathcal{Q}\in\mathbb{R}^{n}$ satisfy $d_{B}\left(\mathcal{S},\mathcal{Q}\right)<r\left(\mathcal{S}\right)$, then we have:
$\mathrm{EMD}\left(\mathcal{G}\left(\mathcal{S}\right),\mathcal{G}\left(\mathcal{Q}\right)\right)$ $=$ (4)
$\mathrm{EMD}\left(\left(\mathcal{S}_{\mathcal{X}},\mathcal{S}_{\mathcal{XI}},\mathcal{S}_{\mathcal{E}},\mathcal{S}_{\mathcal{PDD}}\right),\left(\mathcal{Q}_{\mathcal{X}},\mathcal{Q}_{\mathcal{XI}},\mathcal{Q}_{\mathcal{E}},\mathcal{Q}_{\mathcal{PDD}}\right)\right)$ $=$
$\mathrm{EMD}\left(\left(\mathcal{S}_{\mathcal{X}},\mathcal{Q}_{\mathcal{X}}\right)\right)+\mathrm{EMD}\left(\left(\mathcal{S}_{\mathcal{XI}},\mathcal{Q}_{\mathcal{XI}}\right)\right)$ $+$
$\mathrm{EMD}\left(\left(\mathcal{S}_{\mathcal{E}},\mathcal{Q}_{\mathcal{E}}\right)\right)+\mathrm{EMD}\left(\left(\mathcal{S}_{\mathcal{PDD}},\mathcal{Q}_{\mathcal{PDD}}\right)\right).$
@@
Since disturbances only change the positions of atoms and do not alter their types, therefore $\mathrm{EMD}\left(\left(\mathcal{S}_{\mathcal{X}},\mathcal{Q}_{\mathcal{X}}\right)\right)=0$ and $\mathrm{EMD}\left(\left(\mathcal{S}_{\mathcal{XI}},\mathcal{Q}_{\mathcal{XI}}\right)\right)=0$, refer to $\mathit{Appendix}$ F for details. So, we obtain $\mathrm{EMD}\left(\mathcal{G}\left(\mathcal{S}\right),\mathcal{G}\left(\mathcal{Q}\right)\right)=\mathrm{EMD}\left(\left(\mathcal{S}_{\mathcal{E}},\mathcal{Q}_{\mathcal{E}}\right)\right)+\mathrm{EMD}\left(\left(\mathcal{S}_{\mathcal{PDD}},\mathcal{Q}_{\mathcal{PDD}}\right)\right)\leq 2d_{B}\left(\mathcal{S},\mathcal{Q}\right)$. ∎
@@
The bottleneck distance $d_{B}\left(\mathcal{S},\mathcal{Q}\right)<r\left(\mathcal{S}\right)$ is defined as: $d_{B}\left(\mathcal{S},\mathcal{Q}\right)=\inf_{g:\mathcal{S}\to\mathcal{Q}}\sup_{p\in\mathcal{S}}\left|p-g\left(p\right)\right|$ and the envelope radius $r\left(\mathcal{S}\right)$ is the minimum half-distance between any two points in $r\left(\mathcal{S}\right)$. In other words, $r\left(\mathcal{S}\right)$ is the maximum radius of non-overlapping open balls centered at all points in $\mathcal{S}$. This implies that any small perturbation in atomic positions under the $d_{B}$ carstens1999geometrical will lead to minor changes in the distribution of distances between points in the EMD.
@@
Since the EMD between the constructed crystal graphs only relates to Euclidean distance. Euclidean distance itself is continuous, Theorem 1 extends the following fact: for a unit cell structure with two atoms, when the number of neighbors $k=1$, if we perturb at most two points by $\epsilon$, the change in distance between the two points will be at most $2\epsilon$. Extending this to $n$ atomic points with $k$ neighbors, if we perturb at most $n$ points by $\epsilon$, the change in distance between $n$ points will be at most $2nk\epsilon$. This aligns with Definition 5, hence, the constructed WPDD and UPDD multi-edge crystal graph is continuous.
Method Formation Energy Bandgap(OPT) Total Energy Ehull Bandgap(MBJ) Bulk Moduli(Kv) Shear Moduli(Gv)
eV/atom eV eV/atom eV eV GPa GPa
PotNet(2023) 0.0294 0.127 0.032 0.055 0.27 10.06 8.883
CrysMMNet(2023) 0.028 0.128 0.034 0.278 9.625 8.471
CrysDiff (2024) 0.029 0.131 0.034 0.062 0.287 9.875 9.193
Crystalformer(2024) 0.0306 0.128 0.032 0.046 0.274
eComFormer(2024) 0.0284 0.124 0.032 0.044 0.28 10.79 9.826
iComFormer(2024) 0.0272 0.122 0.0288 0.047 0.26 9.617 9.098
UPDDFormer 0.0267 0.122 0.0287 0.0406 0.260 9.456 8.738
WPDDFormer 0.0257 0.119 0.0276 0.0355 0.249 9.224 8.441
Table 1: Comparison between UPDDFormer, WPDDFormer, and other baselines in terms of test MAE on the JARVIS dataset. The best results are shown in bold and the second-best results are shown with underlines.
Method Formation Energy Band Gap Bulk Shear
eV/atom eV log(GPa) log(GPa)
PotNet 0.0188 0.204 0.040 0.065
CrysMMNet 0.0200 0.197 0.038 0.062
Crystalformer 0.0186 0.198 0.0377 0.0689
eComFormer 0.01816 0.202 0.0417 0.0729
iComFormer 0.01826 0.193 0.0380 0.0637
UPDDFormer 0.01696 0.189 0.0370 0.0670
WPDDFormer 0.01604 0.187 0.0323 0.0605
Table 2: Comparison of test MAE between UPDDFormer, WPDDFormer, and other baselines on the Materials Project dataset.
4.5 General Geometric Completeness
Proposition 2.